OpenStudy (igster1112):

How many representative particles are in 1.45 g of a molecular compound with a molar mass of 237 g? What exactly am I converting to and from?

OpenStudy (igster1112):

@TrojanPoem

OpenStudy (trojanpoem):

N = m/M N: number of moles m: mass M: molar mass N(P) => number of particles N(P) = 6.02 x 10^23 * N

OpenStudy (trojanpoem):

The same as what you solved before. Just get the number of moles from the given mass and molar mass. Then find number of particles with Avogadro's constant.

OpenStudy (igster1112):

So in the top left box I would put 1.45g correct sense it's the given?

OpenStudy (igster1112):

@TrojanPoem I got 3.68*10^20. Would my units be molecules because it does give a specific one.

OpenStudy (igster1112):

*doesn't

OpenStudy (trojanpoem):

particle

OpenStudy (igster1112):

*21*

OpenStudy (igster1112):

Ok. So is the answer correct? 3.67*10^21 particles?

OpenStudy (igster1112):

@TrojanPoem

OpenStudy (trojanpoem):

N= 1.45/237 = 6.12 x 10^-3 N(P) = 6.12 * 10^-3 * 6.02 * 10^23 = 3.68 x 10^21 particles

OpenStudy (igster1112):

*3.68*

OpenStudy (trojanpoem):

Correct bro.

OpenStudy (igster1112):

This one was mass👉moles👉particles right?

OpenStudy (igster1112):

@TrojanPoem

OpenStudy (trojanpoem):

?

OpenStudy (igster1112):

Going from mass to moles to particles?

OpenStudy (trojanpoem):

yeah.

OpenStudy (igster1112):

Ok thank you. That's all I need.

OpenStudy (trojanpoem):

You're welcome